A vector has an x-component of -25.0 units and a y-component of 40.0 units. Find the magnitude and direction of the vector.
1
Expert's answer
2020-01-27T09:46:28-0500

Given:


"a_x=-25.0,\quad a_y=40.0"

The magnitude of the vector "\bf a"


"a=\sqrt{a_x^2+a_y^2}=\sqrt{(-25.0)^2+40.0^2}=47.2"

The angle between vector "\bf a" and positive direction of x-axis


"\theta=\tan^{-1}\frac{a_y}{a_x}=\pi-\tan^{-1}\frac{40.0}{25.0}=122^{\circ}"